# Geometric QCD II: The Confining Twistor String and Meson Spectrum

### 12 Mar 2026 v3

@@ 3 Hodge-dual minimal surface
Now, we are going to define the solvable functional $S[C]$ satisfying both the duality and additivity. We are following the definitions and derivations from [36], adapting them for the present purpose of a base for the fermions living on that minimal surface.
@@ 3.1 The external and internal geometry
We define the surface coordinates $X^{i}_{\mu}(\xi)$ (for $\xi=(\xi_{1},\xi_{2}),i=1,2,3$.  [^2] This surface was first proposed in [33] as a zero-mode of the loop operator in the context of the Yang–Mills gradient flow, but unfortunately, the wrong number of components (four instead of three) was chosen for the $X^{A}_{\mu},A=1,2,3,4$. The rest of the theory in that paper was validated by our recent work [36], including the loop calculus and the Hodge-duality of the minimal surface. The field $X^{i}_{\mu}(\xi)$ maps the disk $\Sigma$ (Fig 2) to $\mathbb{R}^{3}\otimes\mathbb{R}^{4}$.
@@
Note the change of sign, compared with the ordinary area derivative of the minimal surface in four Euclidean dimensions.
Regardless of this sign change, the Hodge duality of the area element $\Sigma$ with these boundary conditions for a generic field $X^{a}_{\mu}(z,\bar{z})$ leads to the same Hodge duality of the area derivative.
This self-duality of the area derivative makes our Hodge-dual minimal area a zero mode of the loop diffusion operator, leading to the solution of the MM loop equations [25, 27], as we argue in the next section.
Note that this area functional is invariant under the reparametrizations of the boundary loop $\delta_{\text{param}}C_{\mu}(\theta)=\epsilon(\theta)\partial_{\theta}C_{\mu}(\theta)$
$\delta_{\text{param}}S=-2\int d\theta\epsilon(\theta)\frac{\partial_{\theta}C_{\mu}\Sigma_{\mu\nu}\partial_{\theta}C_{\nu}}{\sqrt{\Sigma^{2}}}=0$ (3.10)
by the skew symmetry of $\Sigma_{\mu\nu}$.
The Hodge-dual area is defined as a minimum of this functional under an extra duality constraint
$\ast\Sigma_{\mu\nu}\equiv\frac{1}{2}\epsilon_{\mu\nu\alpha\beta}\Sigma_{\alpha\beta}=\chi\Sigma_{\mu\nu}$ (3.11)
In the previous papers [33, 36], we found a complete solution for the Hodge-dual surface in terms of holomorphic map.
@@ 3.2 The Holomorphic Map
So, we have a line in four dimensional complex space $f\in\mathbb{C}^{4}$. Our surface area (3.1) reduces to the induced metric in this space
@@
This area element is Hodge-dual
$\ast\Sigma_{\mu\nu}=\chi\Sigma_{\mu\nu}$ (3.20)
for arbitrary functions $f_{\mu}(z)$.
This duality condition, as argued above, together with the above boundary conditions, provides the self-duality of the area derivative of the minimal area. The self-duality of the area derivative follows from the general formula for variation of the surface $X^{a}_{\mu}(\xi)$ with area (3.1), projected on the holomorphic Ansatz.
@@ 3.3 The Virasoro Constraint and Uniformization
The imposition of the null constraint $(f^{\prime})^{2}=0$ is not an approximation but a rigorous gauge fixing procedure rooted in the geometric symmetries of the problem.
The physical starting point is the geometric Area functional, $S=\int\sqrt{-\det g}\,d^{2}z$, which possesses full reparametrization invariance under general diffeomorphisms $z\to w(z,\bar{z})$. This infinite symmetry group allows us to choose a specific coordinate system on the world-sheet without altering physical results.
We choose to work in isothermal (conformal) coordinates, where the induced metric is diagonal:
$ds^{2}=\rho(z,\bar{z})\,dz\,d\bar{z}\implies g_{zz}=\partial_{z}Y\cdot\partial_{z}Y=0$ (3.22)
For our specific ansatz (3.12), the metric component $g_{zz}$ is explicitly proportional to the square of the holomorphic derivative:
$g_{zz}=(f^{\prime}_{\mu})^{2}=0$ (3.23)
Thus, the geometric condition of conformal gauge fixing $g_{zz}=0$ is algebraically equivalent to the null constraint $(f^{\prime})^{2}=0$.
The existence of such a coordinate system for any surface topology is guaranteed by the Uniformization Theorem [39]. This theorem ensures that we are always free to set $(f^{\prime})^{2}=0$ by a suitable diffeomorphism. In this gauge, the non-linear geometric area density simplifies exactly to the quadratic Lagrangian:
$\mathcal{L}$ $=\sqrt{-\det g}=\sqrt{(g_{z\bar{z}})^{2}-|g_{zz}|^{2}}$
$\xrightarrow{(f^{\prime})^{2}=0}g_{z\bar{z}}\propto|f^{\prime}|^{2}$ (3.24)
This reduction allows us to describe the Minimal Surface using the linear Laplace equation for $f_{\mu}$, subject to the Virasoro constraint. To summarize, we have found that
@@
The solution of the linear boundary problem (3.15) for a holomorphic vector function is given by the Hilbert transform
@@
In terms of the Taylor coefficients
@@
In the presence of the Virasoro constraint (3.21), the problem becomes a challenging one and has been investigated over the last two centuries by great mathematicians, starting with Riemann and Hilbert. We give a brief summary of modern state of this theory in B.
In Part I [36], we compared the ordinary minimal surface in four dimensions with the Hodge-dual one. This comparison reveals some hidden variables present in the area derivatives of the Hodge dual surface, but not in the extremal value of this area. This minimal value up to a normalization factor $2\sqrt{2}$ equals that of the so called Goldschmidt solution [42, 6, 15] for the Plateau problem. This additive solution represents a local minimum of the Dirichlet functional, with the simplest topology (collection of disks one for each closed part of $C$).
The area derivatives for the Goldschmidt minimal area and Hodge dual minimal area are different (the second one being SD or ASD component of the first). The apparent paradox is resolved by the obvious fact that the area derivative of a functional equals the skew-symmetric part of its second functional derivative, but only the extremal value and its first derivative (=0=0) coincide for the functional and its minimal value.
Also, the Hodge-dual surface is not parity-symmetric, but its minimal value (minimal area) is. The chirality $\chi$ manifests itself in the external geometry of this minimal surface $X^{i}_{\mu}\in\mathbb{R}^{3}\otimes\mathbb{R}^{4}$, including its area derivative, but its internal geometry is determined by the induced metric (3.22),(3.23), which is the only local function related to surface embedding. Moreover, this metric in the holomorphic map that minimizes the area depends only on the bounding loop $C$, through the holomorphic vector function $f_{\mu}(z)$, which solves the Riemann-Hilbert problem (9.4), (3.30). This dependence is identical for the Hodge-dual surface and the Goldschmidt solution for the Plateau problem in 4D, bounded by $C$.
The last property (the identity of the internal geometry in the Hodge-dual surface in relation to that of the 4D Euclidean problem) will play an important role in proving the MM equation for the fermionic determinant. The only remnants of the Hodge-duality are the overall negative sign and the $\chi$-term in the area derivative (3.26) projecting it on the Hodge-dual sector of skew-symmetric tensors in 4D.
@@ 4 The Elfin theory on a flat surface
In this section, we deal with the Elfin theory by itself, without any reference to the gauge theory. The equivalence of these theories will be established in the next sections.
@@
This does not invalidate the analysis of the Elfin determinant in the flat gauge: closed fermion paths are weighted by $\exp\left(\!\right)\big(-m\int_{\Gamma}\sqrt{e}\,|dz|\big)=\exp\left((\right)-m\,\ell_{\Gamma})$, hence any path approaching a pole (where $\sqrt{e}\sim|z-w|^{-1}$ for a simple pole of $f^{\prime}$) has $\ell_{\Gamma}\to\infty$ and is exponentially suppressed. Since $\ell_{\Gamma}$ is a local additive functional, the length weights of the locally paired configurations (touching vs. intersecting) coincide, while their mod-$2$ intersection signs are opposite, yielding the same cancelation mechanism.
@@
Let us now discuss the symmetry of the physical $x$-space. As usual in string theory, the transformations of the internal coordinates $\xi$ have nothing to do with the physical rotations in $x$-space. From the viewpoint of two-dimensional field theory in $\xi$-space, the rotations in $x$-space represent internal $O(4)$ symmetry transformations of the $C_{\mu}$ field,
@@
Using the exact area derivatives of the Hodge-dual minimal area established in the previous section, we find for the area derivative of
@@ 5 Conformal metric as a local gauge parameter
is invariant under local variations of the conformal metric field $\rho$. That is, the theory depends only on the conformal class (moduli) of the surface, and the induced metric $e(\xi)$ but not on the local gauge parameter $\rho(\xi)$.
@@ 6.1 Mathematical justification of the signed sum over loops
The path integral representation of the Dirac determinant on a finite surface with boundary conditions, as employed in this work, is rigorously justified by several mathematical results and earlier physical constructions. In particular, Ichinose [19] and Gaveau [13] provide theorems establishing the equivalence between the determinant of the Dirac operator and path integrals over spinor-valued paths with appropriate boundary conditions on finite domains. These constructions carefully incorporate the spinor connection and holonomies of spin rotations along geodesic segments, ensuring the correct treatment of spinor phases and curvature effects.
Our representation of the determinant as a sum over oriented closed loops weighted by factors of the form $\prod\exp\left(-m\,dl+\frac{\mathrm{i}\sigma_{3}\delta\theta}{2}\right)$, where $dl$ is an infinitesimal geodesic step and $\delta\theta$ the corresponding trajectory rotation angle, follows directly from these rigorous path integral formulations. This approach generalizes the combinatorial and topological methods pioneered by Vdovichenko [48] in the exact solution of the 2D Ising model, where the fermionic nature and spinor holonomies enforce cancelations of non-planar intersections, leaving only planar loop configurations that contribute to the determinant.
@@
Thus, the path integral representation employed here is not only physically intuitive but also mathematically sound, grounded in well-established results in spin geometry, spectral theory of Dirac operators, and rigorous path integral constructions on manifolds with boundary. This justifies its use as the foundation for deriving loop equations and analyzing the planar topology of the fermion determinants in the Elfin theory.
@@ 6.3 The Gauss-Bonnet theorem and self-intersections
Mathematical theorems. For an oriented regular simple closed curve $\gamma\subset\mathbb{R}^{2}$, the rotation (turning) index satisfies $\operatorname{rot}(\gamma)=\frac{1}{2\pi}\int_{\gamma}\kappa\,ds\in\mathbb{Z}$ and equals +1+1 for the positively oriented boundary of a disk [53]. In the planar case, this is exactly the Gauss–Bonnet theorem for the enclosed region (since $K\equiv 0$), yielding $\int_{\gamma}\kappa\,ds=2\pi\,\chi=2\pi$ (hence equality modulo $2\pi$ is automatic) [2]. Moreover, since $\operatorname{rot}(\gamma)$ is the degree of the unit-tangent map $T:S^{1}\to S^{1}$, it depends only on the immersed curve itself and therefore is unchanged if the ambient plane is replaced by a disk with finitely many holes (and the same mod-$2$ intersection sign $(-1)^{\nu}$ follows from Whitney’s regular-homotopy classification of immersed plane curves) [53].
@@ 6.4 From exponential of free elfin loops to sum over multiple crossing paths
These loops are independent by construction, but we are going to rearrange the sum in such a way that they will start to repel. This miraculous transformation, first conjectured in [26] is possible due to the following topological and combinatorial observations.
@@ 6.5 Cancellation of intersections: step towards planar graphs
This statement is not as trivial as it seems. The length of the line is not just a geometric length of this line drawn on a planar disk; rather, it is an integral
@@
The non-planar vertex (the second one on the lower drawing in Fig. 3 ) cancels the first one, so this configuration of lines drops from the sum. This is the heart of the matter. The lines may touch each other with opposite orientations according to Fig. 3, but the terms with intersecting lines are always canceled by those with parallel touching lines in the graph with the same configuration of lines.
@@
The time slice ($\xi_{2}=\textit{ const }{}$) of this picture describes a string with particles moving between neighbors. Pairs of $\psi$ and $\bar{\psi}$ particles can be created or annihilated. Only $\psi\bar{\psi}$ neighbors can touch each other. The touching of $\psi\psi$ or $\bar{\psi}\bar{\psi}$ neighbors is forbidden by the Pauli principle. One $\psi,\bar{\psi}$ pair is placed at the two ends of the string.  [^3] This interpretation of the time slice dynamics can be used for a rigorous proof of the equivalence of the free Dirac vacuum loops and the planar loops made of non-intersecting closed paths on a surface. Such a proof would use the Hamiltonian representation of a Fermionic system with conventional creation/annihilation operators and chiral boundary conditions.
@@ 7.1 Loop equations
Eq. (7.1) was postulated in the first version of the Fermi String theory [26] from heuristic arguments. At that time, it seemed that the equation applied to the Ising model at the surface, i.e., for the 2-component spinors. However, as we see now, one should introduce two spinors with special boundary conditions in order to obtain this equation for the propagator of the Fermi string.
@@ 7.2 Large fermion mass and induced QCD
For any minimal surface embedded in Euclidean space, the geodesic distance $L_{geo}(x,y)$ is strictly bounded by the Euclidean distance:
@@ 7.4 The $B$- term
As for the rest of the factors, the sum over the short paths $\Gamma_{up},\Gamma_{dn}$ is dominated by the geodesic distance, which is bounded by the Euclidean distance. This precise alignment relies on the isotropic scaling of White’s bridge in the varifold limit $\epsilon\to 0$. In the strict limit of the topology change, the continuity of the area Hessian demands that the tangent planes perfectly align, yielding the required Fierz-like identity for the MM loop equation without generating anomalous contact terms. A rigorous geometric measure-theory proof of the Hessian’s continuity across this topology change is left for future mathematical study, but is physically mandated by the $O(4)$ symmetry of the local loop diffusion. We find:

### 23 Mar 2026 v4

@@ 3 Hodge-dual minimal surface
We now define the functional $S[C]$ that satisfies the requirements of both duality and additivity. We follow the definitions and derivations from [32], adapting them for the present purpose of a base for the fermions living on that minimal surface.
@@ 3.1 The external and internal geometry
We define the surface coordinates $X^{i}_{\mu}(\xi)$ (for $\xi=(\xi_{1},\xi_{2}),i=1,2,3$. The field $X^{i}_{\mu}(\xi)$ maps the disk $\Sigma$ to $\mathbb{R}^{3}\otimes\mathbb{R}^{4}$.
@@
Note the change of sign, compared with the ordinary area derivative of the minimal surface in four Euclidean dimensions. Regardless of this sign change, the Hodge duality of the area element $\Sigma$ with these boundary conditions for a generic field $X^{a}_{\mu}(z,\bar{z})$ leads to the same Hodge duality of the area derivative. This self-duality of the area derivative makes our Hodge-dual minimal area a zero mode of the loop diffusion operator, leading to the solution of the MM loop equations. The area functional is invariant under the reparametrizations of the boundary loop $\delta_{\text{param}}S=0$ by the skew symmetry of $\Sigma_{\mu\nu}$.
@@ 3.2 The Holomorphic Map
Our surface area (3.1) reduces to the induced metric in this space
@@
This area element is Hodge-dual $\ast\Sigma_{\mu\nu}=\chi\Sigma_{\mu\nu}$ for arbitrary functions $f_{\mu}(z)$. This duality condition provides the self-duality of the area derivative of the minimal area.
@@ 3.3 The Virasoro Constraint and Uniformization
The imposition of the null constraint $(f^{\prime})^{2}=0$ is a gauge fixing procedure rooted in the geometric symmetries of the problem. This allows us to work in isothermal (conformal) coordinates, where the induced metric is diagonal: $g_{zz}=(f^{\prime}_{\mu})^{2}=0$. The existence of such a coordinate system is guaranteed by the Uniformization Theorem. In this gauge, the non-linear geometric area density simplifies to the quadratic Lagrangian:
@@
The solution of the linear boundary problem (3.13) for a holomorphic vector function is given by the Hilbert transform:
@@
In terms of the Taylor coefficients:
@@
The internal geometry of the Hodge-dual surface is identical for the Hodge-dual surface and the Goldschmidt solution for the Plateau problem in 4D. This property is central for establishing the MM equation for the fermionic determinant.
The four-dimensional solution for the minimal surface uses twistors (see B). This is the solution that we use later to build the Regge trajectories.
3.4 The Physical Vacuum, Master field and Multi-Instanton Resummation
Before formulating the dynamics of the Fermi string on the minimal surface, we clarify the origin of the rigid geometric background. The MM loop equations serve as quantum equations of motion for Planar QCD. As with any equations of motion, they admit a family of solutions distinguished by the choice of vacuum. The freedom to multiply the perturbative loop functional by a zero mode reflects this vacuum ambiguity.
In the traditional gauge-field picture, the nonperturbative QCD vacuum is a complex mixed state, often modeled as a distribution in multi-instanton moduli space. Because any individual multi-instanton configuration breaks translation invariance, restoring Poincaré symmetry requires integration over the full moduli space.
In the loop-space formulation, this ensemble is represented geometrically. Just as instantons are characterized by the self-duality of the gauge field strength, the corresponding zero mode is characterized by the self-duality of the loop-space area derivative. The instanton moduli are mapped to the moduli of the twistors parametrizing the four-dimensional minimal surface.
The physical meaning of this geometric representation is analogous to the Master Field proposed by Witten [46], which describes the large-$N_{c}$ limit in terms of a single classical configuration in an infinite-dimensional space. In the present formulation, the Hodge-dual minimal surface plays this role. It provides a classical, translation-invariant object encoding the nonperturbative vacuum, without requiring an explicit construction in terms of gauge-field configurations. In this sense, the minimal surface represents the loop-space realization of the Master Field, with its twistor moduli corresponding to the degrees of freedom associated with the multi-instanton ensemble.
Moreover, the whole surface $X^{i}_{\mu}(\xi)\in\mathbb{R}^{3}\otimes\mathbb{R}^{4}$ has the same algebraic structure as a gauge field in an $SU(2)$ subalgebra. The observables in our theory are the loop functionals and their area derivatives, related to the left and right Gauss maps of the surface. The twistor construction of the minimal surface is thus analogous to a multi-instanton configuration.
Rotations in the internal $\mathbb{R}^{3}$ act as $SO(3)$ transformations on $X^{i}_{\mu}$, corresponding to $SU(2)$ gauge transformations, and physical observables depend only on $SO(3)$-invariant combinations.
The minimal surface described by these twistors is non-fluctuating, reflecting the classical nature of the Master Field in the planar limit.
Thus, the complex mixed state of multi-instantons in coordinate space is mapped, in loop space, to a single translation-invariant geometric object: a Hodge-dual minimal surface. In this representation, the detailed distribution of topological charge and the explicit multi-instanton moduli are not required. The loop equations can be solved directly, without specifying the underlying measure in coordinate space.
The resulting geometric object defines a translation-invariant nonperturbative vacuum for the worldsheet fermions. The degrees of freedom associated with instanton moduli are encoded in the moduli of the minimal surface. In four dimensions, these moduli admit a natural description in terms of holomorphic twistor data.
@@ 4 The Elfin theory on a flat surface
The elfin theory is defined here independently of the gauge theory; the equivalence is established in subsequent sections.
@@
This does not invalidate the analysis of the Elfin determinant in the flat gauge: closed fermion paths are weighted by $\exp\left(-m\int_{\Gamma}\sqrt{e}\,|dz|\right)=\exp\left(-m\,\ell_{\Gamma}\right)$, hence any path approaching a pole (where $\sqrt{e}\sim|z-w|^{-1}$ for a simple pole of $f^{\prime}$) has $\ell_{\Gamma}\to\infty$ and is exponentially suppressed. Since $\ell_{\Gamma}$ is a local additive functional, the length weights of the locally paired configurations (touching vs. intersecting) coincide, while their mod-$2$ intersection signs are opposite, yielding the same cancelation mechanism.
@@
Let us now discuss the symmetry of the physical $x$-space. Transformations of the internal coordinates $\xi$ are independent of physical rotations in $x$-space. From the viewpoint of two-dimensional field theory in $\xi$-space, the rotations in $x$-space represent internal $O(4)$ symmetry transformations of the $C_{\mu}$ field,
@@
Using the area derivatives of the Hodge-dual minimal area established in the previous section, we find for the area derivative of
@@ 5 Conformal metric as a local gauge parameter
is invariant under local variations of the conformal metric field $\rho$. That is, the theory depends only on the conformal class (moduli) of the surface, and the induced metric $e(\xi)$, remaining independent on the local gauge parameter $\rho(\xi)$.
@@ 6.1 Mathematical justification of the signed sum over loops
The path integral representation of the Dirac determinant on a finite surface with boundary conditions, as employed in this work, is justified by several mathematical results and earlier physical constructions. In particular, Ichinose [14] and Gaveau [9] provide theorems establishing the equivalence between the determinant of the Dirac operator and path integrals over spinor-valued paths with appropriate boundary conditions on finite domains. These constructions carefully incorporate the spinor connection and holonomies of spin rotations along geodesic segments, ensuring the correct treatment of spinor phases and curvature effects.
Our representation of the determinant as a sum over oriented closed loops weighted by factors of the form $\prod\exp\left(-m\,dl+\frac{\mathrm{i}\sigma_{3}\delta\theta}{2}\right)$, where $dl$ is an infinitesimal geodesic step and $\delta\theta$ the corresponding trajectory rotation angle, follows directly from these path integral formulations. This approach generalizes the combinatorial and topological methods pioneered by Vdovichenko [41] in the exact solution of the 2D Ising model, where the fermionic nature and spinor holonomies enforce cancelations of non-planar intersections, leaving only planar loop configurations that contribute to the determinant.
@@
Thus, the path integral representation employed here is not only physically intuitive but also mathematically sound, grounded in well-established results in spin geometry, spectral theory of Dirac operators, and path integral constructions on manifolds with boundary. This justifies its use as the foundation for deriving loop equations and analyzing the planar topology of the fermion determinants in the Elfin theory.
@@ 6.3 The Gauss-Bonnet theorem and self-intersections
Mathematical theorems. For an oriented regular simple closed curve $\gamma\subset\mathbb{R}^{2}$, the rotation (turning) index satisfies $\operatorname{rot}(\gamma)=\frac{1}{2\pi}\int_{\gamma}\kappa\,ds\in\mathbb{Z}$ and equals +1+1 for the positively oriented boundary of a disk [45]. In the planar case, this is precisely the Gauss–Bonnet theorem for the enclosed region (since $K\equiv 0$), yielding $\int_{\gamma}\kappa\,ds=2\pi\,\chi=2\pi$ (hence equality modulo $2\pi$ is automatic) [2]. Moreover, since $\operatorname{rot}(\gamma)$ is the degree of the unit-tangent map $T:S^{1}\to S^{1}$, it depends only on the immersed curve itself and therefore is unchanged if the ambient plane is replaced by a disk with finitely many holes (and the same mod-$2$ intersection sign $(-1)^{\nu}$ follows from Whitney’s regular-homotopy classification of immersed plane curves) [45].
@@ 6.4 From exponential of free elfin loops to sum over multiple crossing paths
These loops are independent by construction, but we are going to rearrange the sum in such a way that they will start to repel. This rearrangement, conjectured in [22] relies on the following topological and combinatorial observations.
@@ 6.5 Cancellation of intersections: step towards planar graphs
The length of the line is not just a geometric length of this line drawn on a planar disk; rather, it is an integral
@@
The non-planar vertex (Fig. 3) cancels the first one, removing this configuration from the sum. This is the central mechanism. The lines may touch each other with opposite orientations according to Fig. 3, but the terms with intersecting lines are always canceled by those with parallel touching lines in the graph with the same configuration of lines.
@@
The time slice ($\xi_{2}=\textit{ const }{}$) of this picture describes a string with particles moving between neighbors. Pairs of $\psi$ and $\bar{\psi}$ particles can be created or annihilated. Only $\psi\bar{\psi}$ neighbors can touch each other. The touching of $\psi\psi$ or $\bar{\psi}\bar{\psi}$ neighbors is forbidden by the Pauli principle. One $\psi,\bar{\psi}$ pair is placed at the two ends of the string.  [^2] This interpretation of the time slice dynamics can be used for a proof of the equivalence of the free Dirac vacuum loops and the planar loops made of non-intersecting closed paths on a surface. Such a proof would use the Hamiltonian representation of a Fermionic system with conventional creation/annihilation operators and chiral boundary conditions.
@@ 7.1 Loop equations
Eq. (7.1) was originally proposed in [22]. At that time, it seemed that the equation applied to the Ising model at the surface, i.e., for the 2-component spinors. However, as we see now, one should introduce two spinors with special boundary conditions in order to obtain this equation for the propagator of the Fermi string.
@@ 7.2 Large fermion mass and induced QCD
For any minimal surface embedded in Euclidean space, the geodesic distance $L_{geo}(x,y)$ is bounded by the Euclidean distance:
@@ 7.4 The $B$- term
As for the rest of the factors, the sum over the short paths $\Gamma_{up},\Gamma_{dn}$ is dominated by the geodesic distance, which is bounded by the Euclidean distance. This precise alignment relies on the isotropic scaling of White’s bridge in the varifold limit $\epsilon\to 0$. In the strict limit of the topology change, the continuity of the area Hessian demands that the tangent planes align, yielding the required Fierz-like identity for the MM loop equation without generating anomalous contact terms. A rigorous geometric measure-theory proof of the Hessian’s continuity across this topology change is left for future mathematical study, but is physically mandated by the $O(4)$ symmetry of the local loop diffusion. We find: